Two cars A and B cross a point with velocities 10m/s and 15m/s. After that they move with different uniform accelerations and the car A overtakes B with speed 25m/s. What is the velocity of car B a that instant?
 
[Ans : 20 m/s]

Grade:12

1 Answers

Arun
6 years ago
Dear Ankush
 
Let accn of bodies be a and b respectively. Time after which car a overtakes is t. So we following equations :10 + at = 2510t + 1/2 at^2 = 15t + 1/2 bt^2Solving, at = 15 and bt = 5, so velocity of car b = 15 + bt = 20 m/s
 
Regards
Arun (askIITians forum expert)